OpenStudy (anonymous):

What is the standard equation of this hyperbola? 4x^2 = 36 + 9y^2

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

\[\frac{ x^2 }{ 3^2 }-\frac{ y^2 }{ 5^2}=1\]

OpenStudy (anonymous):

Thats how a is set up

OpenStudy (michele_laino):

we can write this: \[\Large 4{x^2} - 9{y^2} = 36\]

OpenStudy (michele_laino):

now I divide both sides by 36, so I get: \[\Large \frac{{4{x^2}}}{{36}} - \frac{{9{y^2}}}{{36}} = 1\]

OpenStudy (michele_laino):

please simplify the last equation

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

\[6 and 4\]

OpenStudy (anonymous):

\[\frac{ x^2 }{ 4^2 }-\frac{ y^2 }{ 6^2 }=1\]

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: \[\Large \frac{{4{x^2}}}{{36}} - \frac{{9{y^2}}}{{36}} = \frac{{{x^2}}}{{\frac{{36}}{4}}} - \frac{{{y^2}}}{{\frac{{36}}{9}}}\]

OpenStudy (anonymous):

so its 2^2 and 3^2

OpenStudy (michele_laino):

yes! it is: \[\Large \frac{{4{x^2}}}{{36}} - \frac{{9{y^2}}}{{36}} = \frac{{{x^2}}}{{\frac{{36}}{4}}} - \frac{{{y^2}}}{{\frac{{36}}{9}}} = \frac{{{x^2}}}{{{3^2}}} - \frac{{{y^2}}}{{{2^2}}} = 1\]